One base angle of a right triangle has the measure of 28 degrees find the measure of the other acute angle. How would i do this?
1
Expert's answer
2013-01-10T08:04:19-0500
It's known that sum of the angles of every triangle equals 180 degrees. As considered triangle is right, it has right angle measuring 90 degrees. Also we're given measure of one of its acute angles - 28 degrees. Let measure of unknown angle is x. Then we can construct the following equation:
90+28+x=180
whence x=180-90-28=62